### v1

# On the structure of the Birkhoff-irregular set for some Subshifts of Finite Type

Abstract.
We study the set of irregular points for a class of topologically mixing subshifts of finite type. It is well known that despite the irregular set having zero measure for every invariant measure, it has full topological entropy and full Hausdorff dimension. We establish that for these systems the irregular set is not only abundant in terms of its dimensional properties, but also contains uncountably many pairwise disjoint subsets, each with full entropy and dimension. This result deepens our understanding of the complexity of irregular points in dynamical systems, highlighting their intricate structure and suggesting avenues for further exploration in related areas.
1. Introduction
Given a compact metric space $(X,d)$ and a continuous function $f\colon X\to X$, a point $x_{0}\in X$ is called (Birkhoff) irregular if there is a continuous function $\varphi\colon X\to\mathbb{R}$ such that its Birkhoff average at $x_{0}$
$\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\varphi(f^{j}(x_{0}))$
does not exist. The irregular set of $\varphi$ is denoted by $\hat{X}(\varphi)=\hat{X}(f,\varphi)$ and the irregular set of $X$ is denoted by $\hat{X}=\hat{X}(f)$. It is well known by Birkhoff ergodic theorem that for every $f$-invariant measure $\mu$ we have $\mu(\hat{X})=0$, so the irregular set is “invisible” from the measure-theoretic viewpoint. Thus, it is natural to measure its size using other functions. We will focus on the most common approaches, one of them geometric in nature, using Hausdorff dimension, and another dynamic in nature, using topological entropy. For topological entropy, we use its Carathéodory dimensional definition, where it is not required that the set is compact or invariant.
Pesin and Pitskel [PesinPitskel] studied the irregular set for the full shift on two symbols, and showed that it has full topological entropy. Barreira and Schmeling [BS] showed that for subshifts of finite type, the irregular set carries full topological entropy and full Hausdorff dimension, and applied these results to hyperbolic dynamical systems. Thompson [T1] studied the irregular set for any topological dynamical system satisfying the specification property, and showed that it has full topological pressure. Later Thompson [T2] showed that the irregular set for systems with the almost specification property has full topological entropy and full Hausdorff dimension. Gelfert, Pacifico and Sanhueza [GPS] gave sufficient conditions for a dynamical system to have irregular set with full topological entropy, applied to examples including nonuniformly hyperbolic diffeomorphisms with positive entropy and rational functions on the Riemann sphere. These results give us the property that even though the irregular set is negligible from the viewpoint of invariant measures, many times it is expected to have the largest possible size in terms of Hausdorff dimension and entropy.
We focus on the structure of the irregular set in the following sense: how many pairwise disjoint sets of full dimension and entropy can we have inside the irregular set? How are these sets distributed inside $X$? We answer these questions in the following theorem.
Theorem 1.1.
Let $(X,\sigma)$ be a topologically mixing subshift of finite type. Assume that there are two symbols that can be preceeded and succeeded by any other symbol. Then, there is an uncountable collection $\mathcal{N}$ of subsets of $X$ satisfying:
  1. (1)
    There is $\varphi\in C(X)$ such that every $N\in\mathcal{N}$ is contained in $\hat{X}(\varphi)$;
  2. (2)
    every $N\in\mathcal{N}$ is dense in $X$ and satisfies $\sigma(N)\subset N$;
  3. (3)
    for every distinct $N,N^{\prime}\in\mathcal{N}$ we have $N\cap N^{\prime}=\emptyset$;
  4. (4)
    for every $N\in\mathcal{N}$ we have $\dim_{H}N=\dim_{H}X$ and $h(\sigma|N)=h_{top}(\sigma)$.
Our work follows ideas from Pesin and Pitskel [PesinPitskel], which are adapted to some subshifts of finite type. We will work on one-sided subshifts of finite type, though all the arguments work as well in the case of two-sided subshifts. Since horseshoes and iterated functions systems can be coded using full shifts, and conjugacies preserve entropy, we obtain the following corollaries as direct applications of Theorem 1.1.
Corollary 1.2.
Let $f\colon\Lambda\to\Lambda$ be a Smale horseshoe map. Then, there is an uncountable collection $\mathcal{N}$ of subsets of $\Lambda$ satisfying:
  1. (1)
    There is $\varphi\in C(\Lambda)$ such that every $N\in\mathcal{N}$ is contained in $\hat{\Lambda}(\varphi)$;
  2. (2)
    every $N\in\mathcal{N}$ is dense in $\Lambda$ and satisfies $\sigma(N)\subset N$;
  3. (3)
    for every distinct $N,N^{\prime}\in\mathcal{N}$ we have $N\cap N^{\prime}=\emptyset$;
  4. (4)
    for every $N\in\mathcal{N}$ we have $h(f|N)=h_{top}(f)$.
Corollary 1.3.
Let $X\subset\mathbb{R}^{d}$ be the attractor of an iterated function system $(T_{1},\ldots,T_{N})$ satisfying the strong separation condition. Consider the dynamics $f\colon X\to X$ defined by the inverses of the maps $T_{i}$. Then, there is an uncountable collection $\mathcal{N}$ of subsets of $X$ satisfying:
  1. (1)
    There is $\varphi\in C(X)$ such that every $N\in\mathcal{N}$ is contained in $\hat{X}(\varphi)$;
  2. (2)
    every $N\in\mathcal{N}$ is dense in $X$ and satisfies $\sigma(N)\subset N$;
  3. (3)
    for every distinct $N,N^{\prime}\in\mathcal{N}$ we have $N\cap N^{\prime}=\emptyset$;
  4. (4)
    for every $N\in\mathcal{N}$ we have $h(f|N)=h_{top}(f)$.
Acknowledgements. I would like to express my sincere appreciation to my advisor, Yakov Pesin, for his invaluable guidance and support throughout this research. I also wish to thank Scott Schmieding for the insightful discussions that contributed to this work. Additionally, I would like to thank Sebastian Pavez, Ygor de Jesus, and the Penn State student dynamics group for their encouragement and comments.
2. Hausdorff dimension and Topological entropy
We follow the definitions of Hausdorff dimension and topological entropy from [missing], using Carathéodory structures.
2.1. Carathéodory Structures
Let $X$ be a set and let $I$ be a set of indices. Let =$\mathcal{F}=\{U_{i}:i\in I\}$ be a collection of subsets of $X$. Assume that there exist functions $\xi,\psi,\eta\colon I\to[0,\infty)$ satisfying:
  1. C1.
    there is $i_{0}\in I$ such that $U_{i_{0}}=\emptyset$; if $U_{i}=\emptyset$ then $\psi(i)=\eta(i)=0$; if $U_{i}\neq\emptyset$ then $\psi(i)>0$ and $\eta(i)>0$;
  2. C2.
    for every $\delta>0$ there is $\varepsilon>0$ such that $\eta(i)\leq\delta$ for any $i\in I$ with $\psi(i)\leq\varepsilon$;
  3. C3.
    for any $\varepsilon>0$ there is a finite or countable collection $J\subset I$ that covers $X$ (i.e. $X\subset\bigcup_{i\in J}U_{j}$) and $\psi(J):=\sup\{\psi(i):i\in J\}\leq\varepsilon$.
We say that the tuple $\tau=(\mathcal{F},\xi,\eta,\psi)$ is a Carathéodory dimension structure (or C-structure).
For $\alpha\in\mathbb{R}$ and $Z\subset X$, the expression
(1) $m_{C}(Z,\alpha):=\lim_{\varepsilon\to 0}\inf\left\{\sum_{i\in J}\xi(i)\eta(i)^{\alpha}:J\subset I\text{ covers }Z,\psi(J)\leq\varepsilon\right\}$
defines the $\alpha$-Caratéodory outer measure on $X$, which induces a $\sigma$-additive measure on $X$. It can be shown [missing, Proposition 1.2] that given $Z\subset X$, there is a critical value $-\infty\leq\alpha_{\star}\leq\infty$ such that $m_{C}(Z,\alpha)=\infty$ for every $\alpha<\alpha_{\star}$ and $m_{C}(Z,\alpha)=0$ for every $\alpha>\alpha_{\star}$.
Definition 2.1.
The Carathéodory dimension of a set $Z\subset X$ is defined by
(2) $\dim_{C}Z:=\inf\{\alpha:m_{C}(Z,\alpha)=0\}=\sup\{\alpha:m_{C}(Z,\alpha)=\infty\}.$
This Carathéodory dimension satisfies the following properties [missing, Theorem 1.1].
Proposition 2.2.
The following hold:
  1. (1)
    $\dim_{C}\emptyset\leq 0$;
  2. (2)
    if $Z_{1}\subset Z_{2}\subset X$, then $\dim_{C}Z_{1}\leq\dim_{C}Z_{2}$;
  3. (3)
    $\dim_{C}\left(\displaystyle\bigcup_{i\in\mathbb{N}}Z_{i}\right)=\displaystyle\sup_{i\in\mathbb{N}}\dim_{C}Z_{i}$.
Now suppose $(X,\mu)$ is a probability space such that $\mathcal{F}$ is contained in the measurable sets of $X$. We define the Carathéodory dimension of $\mu$ by
(3) $\dim_{C}\mu:=\inf\{\dim_{C}Z:\mu(Z)=1\}.$
2.2. Hausdorff Dimension
Let $(X,d)$ be a complete separable metric space and let $\mathcal{F}=I$ be the set of all open subsets of $X$. Consider $\xi,\psi,\eta\colon\mathcal{F}\to[0,\infty)$ defined by
(4) $\xi(U)=1,\qquad\psi(U)=\eta(U)=\operatorname{diam}(U).$
Definition 2.3.
Let $Z\subset X$. The Carathéodory dimension of $Z$ given by the $C$-structure $\tau=(\mathcal{F},\xi,\eta,\psi)$ defined by (4) is called the Hausdorff dimension of $Z$, and is denoted $\dim_{H}Z$.
Assume now that $X$ is a Besicovitch metric space, then we have a way of computing the Hausdorff dimension of measures [missing, Theorem 7.1 and Appendix I]. Given a Borel probability measure $\mu$ on $X$, its pointwise dimension at $x\in X$ is defined by
(5) $d_{\mu}(x):=\lim_{r\to 0}\dfrac{\log\mu(B(x,r))}{\log r},$
whenever the limit exists.
Proposition 2.4.
If $d_{\mu}(x)=d$ for $\mu$-almost every $x\in X$, then $\dim_{H}\mu=d$.
We also have the following proposition as immediate corollary of Proposition 2.2.
Proposition 2.5.
The following hold:
  1. (1)
    $\dim_{H}\emptyset=0$;
  2. (2)
    if $Z_{1}\subset Z_{2}\subset X$, then $\dim_{H}Z_{1}\leq\dim_{H}Z_{2}$;
  3. (3)
    $\dim_{H}\left(\displaystyle\bigcup_{i\in\mathbb{N}}Z_{i}\right)=\displaystyle\sup_{i\in\mathbb{N}}\dim_{H}Z_{i}$.
2.3. Topological entropy
Let $(X,d)$ be a compact metric space and let $f\colon X\to X$ be a continuous map. Let $\mathcal{U}$ be an open cover of $X$ and let
$\mathcal{S}=\mathcal{S}_{m}(\mathcal{U}):=\{{\bf U}=U_{i_{0}}\cdots U_{i_{m-1}}:U_{i_{j}}\in\mathcal{U}\}$
be the set of all strings of length $m$ of elements of $\mathcal{U}$. Denote by $m({\bf U})$ the length of the string ${\bf U}$ and set $\mathcal{S}(\mathcal{U}):=\bigcup_{m\geq 0}\mathcal{S}_{m}(\mathcal{U})$, which will be our set of indices.
To a string ${\bf U}\in\mathcal{S}(\mathcal{U})$ we assign the set
(6) $X({\bf U})=\{x\in X:f^{j}(x)\in U_{i_{j}}\text{ for }j=0,\ldots,m({\bf U})-1\}.$
Consider the collection of subsets $\mathcal{F}=\mathcal{F}(\mathcal{U})=\{X({\bf U}):{\bf U}\in\mathcal{S}(\mathcal{U})\}$ and define functions $\xi,\psi,\eta\colon\mathcal{S}(\mathcal{U})\to[0,\infty)$ by
(7) $\xi({\bf U}):=1,\qquad\eta({\bf U}):=e^{-m({\bf U})},\qquad\psi({\bf U})=\frac{1}{m({\bf U})}.$
These functions determine a $C$-structure $\tau=\tau(\mathcal{U})=(\mathcal{S},\mathcal{F},\xi,\eta,\psi)$. Observe that the Carathéodory outer measure, and hence the Carathéodory dimension, depend on the open cover $\mathcal{U}$. Given $Z\subset X$, we denote its Carathéodory dimension given by this structure by $h_{\mathcal{U}}(f|Z):=\dim_{C,\mathcal{U}}Z$.
Let $\operatorname{diam}(\mathcal{U}):=\sup\{\operatorname{diam}(U):U\in\mathcal{U}\}$ be the diameter of the cover $\mathcal{U}$. Then, Theorem 11.1 in [missing] allows us to make the following definition.
Definition 2.6.
Let $Z\subset X$. The topological entropy of $f$ restricted to the set $Z$ is defined by
(8) $h(f|Z):=\lim_{\operatorname{diam}(\mathcal{U})\to 0}h_{\mathcal{U}}(f|Z).$
We have the corresponding proposition for topological entropy.
Proposition 2.7.
The following hold:
  1. (1)
    $h(f|\emptyset)=0$;
  2. (2)
    if $Z_{1}\subset Z_{2}\subset X$, then $h(f|Z_{1})\leq h(f|Z_{2})$;
  3. (3)
    $h\left(f|\bigcup_{i\in\mathbb{N}}Z_{i}\right)=\displaystyle\sup_{i\in\mathbb{N}}h(f|Z_{i})$.
It is important to mention that for a compact invariant set $Z$, the topological entropy defined using this Carathéodory structure coincide with the classical definition of topological entropy.
3. Subshifts of Finite Type
3.1. Basic definitions
Let $A$ be a $d\times d$ matrix with entries in $\{0,1\}$. The compact set
$X=\{(x_{i})_{i\geq 0}:x_{i}\in\{0,1,\ldots d-1\}\text{ and }A_{x_{i}x_{i+1}}=1\text{ for every }i\}$
is called a topological Markov shift, and $A$ is called its adjacency (or transfer) matrix. The shift map $\sigma\colon X\to X$ is defined by $\sigma((x_{i})_{i\geq 0})=(x_{i+1})_{i\geq 0}$. The dynamical system $(X,\sigma)$ is called a (one-sided) subshift of finite type. For $x=(x_{i})_{i\geq 0}$ and $y=(y_{i})_{i\geq 0}$, define a distance on $X$ by
(9) $d(x,y):=e^{-\min\{i\geq 0:x_{i}\neq y_{i}\}}$
when $x\neq y$, and $d(x,y)=0$ when $x=y$.
The matrix $A$ is called aperiodic if there is $k\in\mathbb{N}$ such that all the entries of $A^{k}$ are positive. This is equivalent to the dynamical system $(X,\sigma)$ being topologically mixing, i.e. for any two open sets $U,V\subset X$, there is $N\in\mathbb{N}$ such that $\sigma^{n}(U)\cap V\neq\emptyset$ for every $n\geq N$.
Let $\mathcal{L}_{n}(X)$ be the set of words in the alphabet $\{0,1,\ldots,d-1\}$ which have length $n$. Given a word $w\in\mathcal{L}_{n}(X)$, the cylinder set centered at $w$ is defined by
$[w]:=\{x\in X:x_{0}=w_{0},\ldots,x_{n-1}=w_{n-1}\}.$
Given $x\in X$, define the cylinder of length $n$ centered at $x$ by $C_{n}(x):=[x_{0}\ldots x_{n-1}]$, which coincide with the open ball centered at $x\in X$ of length $e^{-n}$. These sets generate the topology on $X$ induced by the metric $d$ defined in (9).
The following theorem is well known, and will be very important for us describing the measure of maximal entropy of $(X,\sigma)$.
Theorem 3.1 (Perron-Frobenius Theorem).
Let $A$ be a non-negative aperiodic $d\times d$ matrix. Then, the following hold:
  1. (1)
    There is a positive real eigenvalue $\lambda$ of $A$ such that all other eigenvalues $\rho\in\mathbb{C}$ of $A$ satisfy $\lvert\rho\rvert<\lambda$.
  2. (2)
    The eigenvalue $\lambda$ is simple. That is, the corresponding eigenspace is one-dimensional.
  3. (3)
    There are unique left- and right- eigenvectors ${\bf u}$ and ${\bf v}$ of $A$ corresponding to the eigenvalue $\lambda$ with $u_{i}>0$, $v_{i}>0$ for all $i$, also satisfying $\sum_{i}u_{i}=1$ and ${\bf u}^{T}{\bf v}=1$.
3.2. Markov measures
A $d\times d$ matrix $P$ is called stochastic if its entries are between 0 and 1, and the elements of each of its rows add up to 1. A vector ${\bf p}\in\mathbb{R}^{n}$ is called a probability vector if it has coordinates between 0 and 1, and their sum equals 1. A stochastic matrix $P$ is compatible with $X$ if $P_{ij}>0$ only as long as $A_{ij}>0$. If $P$ is aperiodic, there is a unique probability vector ${\bf p}$ satisfying ${\bf p}^{T}P={\bf p}$.
Definition 3.2.
The Markov measure $\mu_{P}$ associated to $(P,{\bf p})$ is defined in cylinders as follows. For $w\in\mathcal{L}_{n}(X)$, let
(10) $\mu_{P}([w]):=p_{w_{0}}P_{w_{0}w_{1}}P_{w_{1}w_{2}}\cdots P_{w_{n-2}w_{n-1}}.$
Markov measures are ergodic $\sigma$-invariant Borel probability measures on $X$. Moreover, these measures have exponential decay of correlations with respect to Hölder observables (see for example [missing, Theorem 7.4.1]), i.e. there is $\gamma>0$ such that for every $\phi,\psi\colon X\to\mathbb{R}$ Hölder continuous, we have
[…]
for all $n\in\mathbb{N}$, where $C=C(\phi,\psi)$.
The measure-theoretic entropy of a Markov measure $\mu_{P}$ is given by
(11) $h(\mu_{P})=\sum_{i=0}^{d-1}p_{i}\sum_{j=0}^{d-1}-P_{ij}\log P_{ij}.$
3.3. Entropy, dimension and the Parry measure
We will now state some facts about Hausdorff dimension of ergodic measures, as well as entropy and Hausdorf dimension of the entire space $X$. Recall that we are considering the metric defined by (9). We denote by $\mathcal{M}_{\sigma}(X)$ the set of all $\sigma$-invariant probability measures on $X$, and for any $\mu\in\mathcal{M}_{\sigma}(X)$ we denote by $h(\mu)$ its measure theoretical entropy.
The first proposition is an immediate corollary of Shannon-McMillan Breiman theorem and Proposition 2.4. The second proposition can be found in [missing, Theorem A.2.9].
Proposition 3.3.
Let $(X,\sigma)$ be a subshift of finite type. Then, for any ergodic measure $\mu$ on $X$ we have $\dim_{H}\mu=h(\mu)$.
Proposition 3.4.
Let $(X,\sigma)$ be a subshift of finite type with adjacency matrix $A$. Then, $\dim_{H}X=h_{top}(\sigma)=\log\lambda$, where $\lambda$ is as in Theorem 3.1.
The well known Variational Principle states that
$h_{top}(\sigma)=\sup\{h(\mu):\mu\in\mathcal{M}_{\sigma}(X)\},$
where any measure that attains the supremum (if any) is called a measure of maximal entropy of $(X,\sigma)$. Parry [parry] showed that if $(X,\sigma)$ is a topologically mixing subshift of finite type, then there is an aperiodic stochastic matrix $P$ such that $m:=\mu_{P}$ is the unique measure of maximal entropy for $(X,\sigma)$. Such measure is called the Parry measure. The matrix $P$ and its left eigenvector ${\bf p}$ of the Parry measure are explicitly given by
(12) $P_{ij}=A_{ij}\frac{v_{j}}{\lambda v_{i}},\qquad p_{i}=u_{i}v_{i},$
where $\lambda$, ${\bf u}$, ${\bf v}$ are as in Theorem 3.1.
[…]
(13) $m([iwj])=\frac{u_{i}v_{j}}{\lambda^{n}}.$
3.4. Main theorem
In this section we re-state the main result of the paper. Let $(X,\sigma)$ be a topologically mixing subshift of finite type with adjacency matrix $A$ and metric defined by (9).
Definition 3.5.
We call a symbol $a\in\{0,1,\ldots,d-1\}$ safe for $X$ if it can be preceded and succeded by any other symbol. That is, the adjacency matrix satisfies $A_{ai}=A_{ja}=1$ for every $11\leq i,j\leq d-$ 1.
Theorem 3.6.
Assume $X$ has two different safe symbols. Then, there is an uncountable collection $\mathcal{N}$ of subsets of $X$ satisfying:
  1. (1)
    There is $\varphi\in C(X)$ such that every $N\in\mathcal{N}$ is contained in $\hat{X}(\varphi)$;
  2. (2)
    every $N\in\mathcal{N}$ is dense in $X$ and satisfies $\sigma(N)\subset N$;
  3. (3)
    for every distinct $N,N^{\prime}\in\mathcal{N}$ we have $N\cap N^{\prime}=\emptyset$;
  4. (4)
    for every $N\in\mathcal{N}$ we have $\dim_{H}N=h(\sigma|N)=\log\lambda$.
4. Proof of the main theorem
The idea of the proof is to find a sequence of ergodic measures such that their entropies (resp. Hausdorff dimensions) approximate the maximal entropy (resp. Hausdorff dimension). Thus, the union of some of their regular sets will approximate the desired quantities. Following the ideas of [PesinPitskel], we will map these regular sets onto irregular sets without distorting the maximal entropy and dimension. By Proposition 3.3, we can refer to only entropy of the measures, as the Hausdorff dimension coincide.
Without loss of generality, we may assume that the safe symbols are 0 and 1. Fix $P$ and ${\bf p}$ such that the corresponding Markov measure $m=\mu_{P}$ is the Parry measure. In order to approximate entropy of $m$, we will perturb the entries of the matrix
$P=\begin{pmatrix}P_{00}&P_{01}&P_{02}&\cdots&P_{0(d-1)}\\ P_{10}&P_{11}&P_{12}&\cdots&P_{1(d-1)}\\ P_{20}&P_{21}&P_{22}&\cdots&P_{1(d-1)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ P_{(d-1)0}&P_{(d-1)1}&P_{(d-1)2}&\cdots&P_{(d-1)(d-1)}\end{pmatrix}.$
Lemma 4.1.
For every $10\leq i\leq d-$ 1 we have $P_{i0}=P_{i1}$ and $P_{0i}=P_{1i}$.
Proof.
The multiplication $A{\bf v}=\lambda{\bf v}$ gives us
$\sum_{j=0}^{d-1}v_{j}=\lambda v_{1}=\lambda v_{2},$
so $v_{0}=v_{1}$. Now let $10\leq i\leq d-$ 1. Since $A_{i0}=A_{i1}=A_{0i}=A_{1i}=1$, by (12) we have
$P_{i0}=A_{i0}\frac{v_{0}}{\lambda v_{i}}=A_{i1}\frac{v_{1}}{\lambda v_{i}}=P_{i1}\quad\text{and}\quad P_{0i}=A_{0i}\frac{v_{i}}{\lambda v_{0}}=A_{1i}\frac{v_{i}}{\lambda v_{1}}=P_{1i}.\qed$
Consider the matrix
(14) $Q=\begin{pmatrix}\alpha&\gamma&P_{02}&\cdots&P_{0(d-1)}\\ \alpha&\gamma&P_{12}&\cdots&P_{1(d-1)}\\ P_{20}&P_{21}&P_{22}&\cdots&P_{1(d-1)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ P_{(d-1)0}&P_{(d-1)1}&P_{(d-1)2}&\cdots&P_{(d-1)(d-1)}\end{pmatrix},$
where $0<\alpha<P_{00}$ and $\gamma:=1-\alpha-\sum_{j=2}^{d-1}P_{0j}=1-\alpha-\sum_{j=2}^{d-1}P_{1j}$ (by Lemma 4.1). Thus, $Q$ is a stochastic matrix. Since $P$ is aperiodic, there is a power of this matrix with positive entries. By continuity, if $\alpha$ is sufficiently close to $P_{00}$, the matrix $Q$ is also aperiodic, and hence it has a unique left eigenvector ${\bf q}$ with eigenvalue 1 which defines a Markov measure $\mu=\mu_{Q}\in\mathcal{M}_{\sigma}(X)$.
Consider $\chi_{0},\chi_{1}\in C(X)$ defined by
$\chi_{0}(x)=\left\{\begin{array}[]{ccl}1&\text{if}&x_{0}=0\\ 0&\text{if}&x_{0}\neq 0\end{array}\right.\qquad\text{and}\qquad\chi_{1}(x)=\left\{\begin{array}[]{ccl}1&\text{if}&x_{0}=1\\ 0&\text{if}&x_{0}\neq 1\end{array}\right.,$
and let $B(\mu)$ be the set of $x\in X$ such that
$\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\chi_{0}(\sigma^{j}(x))=\mu([0])=q_{0}\quad\text{and}\quad\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\chi_{1}(\sigma^{j}(x))=\mu([1])=q_{1},$
which satisfies $\mu(B(\mu))=1$ by Birkhoff Ergodic Theorem.
Let $S_{1}$ be the subset of natural numbers $i$ such that $(2k)!\leq i\leq(2k+1)!$ for some $k\in\mathbb{N}$ and let $S_{2}:=\mathbb{N}\setminus S_{1}$. Define a function $L\colon X\to X$ as follows: for $x=(x_{i})_{i\geq 0}$,
  • if $i\in S_{1}$, set $L(x)_{i}=x_{i}$;
  • if $i\in S_{2}$, set
    $L(x)_{i}=\left\{\begin{array}[]{ccl}0&\text{if}&x_{i}=1\\ 1&\text{if}&x_{i}=0\\ x_{i}&\text{if}&x_{i}\in\{2,3,\ldots,d-1\}\end{array}\right..$
The map $L$ is well defined as 0 and 1 are safe symbols. Moreover, $L$ is a bijection that preserves distances. We claim that $L(B(\mu))$ consists of irregular points. First, we need the following lemma.
Lemma 4.2.
Let ${\bf q}$ be the left eigenvector of the matrix $Q$ with eigenvalue 1. Then, $q_{0}<q_{1}$.
Proof.
Since ${\bf q}^{T}Q={\bf q}^{T}$, using the fact that $\alpha<\gamma$ and Lemma 12 we have
$q_{0}=q_{0}\alpha+q_{1}\alpha+\sum_{j=2}^{d-1}q_{j}P_{j0}<q_{0}\gamma+q_{1}\gamma+\sum_{j=2}^{d-1}q_{j}P_{j1}=q_{1}.\qed$
Lemma 4.3.
Let $\varphi:=\chi_{0}$. Then, $L(B(\mu))\subset\hat{X}(\varphi)$.
Proof.
Fix $x\in B(\mu)$, and observe that if $(2n-1)!+1\leq j\leq(2n)!-1$ for some $n\geq 0$ we have
$\chi_{0}(\sigma^{j}(L(x)))=\left\{\begin{array}[]{ccl}1&\text{if}&L(x)_{j}=0\\ 0&\text{if}&L(x)_{j}\neq 0\end{array}\right.=\left\{\begin{array}[]{ccl}1&\text{if}&x_{j}=1\\ 0&\text{if}&x_{j}\neq 1\end{array}\right.=\chi_{1}(\sigma^{j}(x)).$
Since
$\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\chi_{0}(\sigma^{j}(x))=q_{0}\quad\text{and}\quad\lim_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}\chi_{1}(\sigma^{j}(x))=q_{1},$
we have
$\Big{|}\frac{1}{(2n+1)!}\sum_{i=0}^{(2n+1)!}\chi_{0}(\sigma^{i}(L(x)))-q_{0}\Big{|}$ $\leq\Big{|}\frac{1}{(2n+1)!}\sum_{i=0}^{(2n+1)!}\chi_{0}(\sigma^{i}(L(x)))-\frac{1}{(2n+1)!}\sum_{i=0}^{(2n+1)!}\chi_{0}(\sigma^{i}(x))\Big{|}$
$+\Big{|}\frac{1}{(2n+1)!}\sum_{i=0}^{(2n+1)!}\chi_{0}(\sigma^{i}(x))-q_{0}\Big{|}$
$\leq\frac{1}{2n+1}+\Big{|}\frac{1}{(2n+1)!}\sum_{i=0}^{(2n+1)!}\chi_{0}(\sigma^{i}(x))-q_{0}\Big{|}\longrightarrow 0$
as $n\to\infty$. Similarly,
$\Big{|}\frac{1}{(2n)!}\sum_{i=1}^{(2n)!-1}\chi_{0}(\sigma^{i}(L(x)))-q_{1}\Big{|}$ $\leq\Big{|}\frac{1}{(2n)!}\sum_{i=1}^{(2n)!-1}\chi_{0}(\sigma^{i}(L(x)))-\frac{1}{(2n)!}\sum_{i=1}^{(2n)!-1}\chi_{1}(\sigma^{i}(x))\Big{|}$
$+\Big{|}\frac{1}{(2n)!}\sum_{i=1}^{(2n)!-1}\chi_{1}(\sigma^{i}(x))-q_{1}\Big{|}$
$\leq\frac{1}{2n}+\Big{|}\frac{1}{(2n)!}\sum_{i=1}^{(2n)!-1}\chi_{1}(\sigma^{i}(x))-q_{1}\Big{|}\longrightarrow 0$
as $n\to\infty$. Since $q_{0}<q_{1}$ (by Lemma 4.2), $L(x)\in\hat{X}(\varphi)$. The desired result follows as $x\in B(\mu)$ was arbitrary.∎
4.1. Constructing a set $N$ with properties (1) and (2) in Theorem 3.6
Consider an increasing sequence $\mathcal{P}=\{\alpha_{k}:k\geq 1\}$ in $(0,P_{00})$ converging to $P_{00}$, such that the matrices $Q_{k}$ defined as in (14) (with $\alpha_{k}$ instead of $\alpha$) are all aperiodic, and therefore they have unique left eigenvectors ${\bf q}_{k}$ with eigenvalue 1. Let $\mu_{k}$ be the Markov measure associated to $(Q_{k},{\bf q}_{k})$.
For every $k\geq 1$ let $B_{k}:=B(\mu_{k})$ be as in the previous subsection (using the 0th and 1st coordinates of the vector ${\bf q}_{k}$), and define $N:=\bigcup_{k\geq 1}L(B_{k})$. By Lemma 4.3, for each $k\in\mathbb{N}$ we have $L(B_{k})\subset\hat{X}(\varphi)$, so $N\subset\hat{X}(\varphi)$.
Lemma 4.4.
The set $N$ is dense in $X$.
Proof.
Since the vectors ${\bf u},{\bf v}$ in Theorem 3.1 have positive entries, by (13) the Parry measure $m$ gives positive measure to any cylinder in $X$.
Let $C$ be an arbitrary cylinder in $X$. By the previous observation, $m(C)>0$. By continuity of the expression (10) depending on the matrix $Q_{k}$ and the vector ${\bf q}_{k}$, there is $k_{0}\in\mathbb{N}$ such that $\mu_{k_{0}}(C)>0$. Thus, $B_{k_{0}}\cap C\neq\emptyset$ as $\mu_{k_{0}}(B_{k_{0}})=1$. We conclude that the set $\bigcup_{k\geq 1}B_{k}$ is dense in $X$. Since $L$ is a homeomorphism of $X$, the set $N=L(\bigcup_{k\geq 1}B_{k})$ is also dense in $X$. ∎
Lemma 4.5.
We have $\sigma(N)\subset N$.
Proof.
Let $y\in N$. Then, there is $k\in\mathbb{N}$ and $x\in B_{k}$ such that $L(x)=y$. We need to show that $\sigma(y)\in N$, so it suffices to show that $L(\sigma(y))\in B_{k}$.
Let $n\in\mathbb{N}$ and let $m(n)$ be the maximum natural number $m$ such that $m!\leq n-1$. Observe that $m(n)\to\infty$ as $n\to\infty$. We have that the first $n$ terms of the sequences $x=(x_{i})_{i\in\mathbb{N}_{0}}$ and $(L\circ\sigma\circ L)(x)$ may disagree on at most $m(n)$ symbols. Then, for $j=1,2$ we have
$\Big{|}\frac{1}{n}\sum_{i=0}^{n-1}\chi_{j}(\sigma^{i}(x))-\frac{1}{n}\sum_{i=0}^{n-1}\chi_{j}(\sigma^{i}((L\circ\sigma\circ L)(x)))\Big{|}\leq\frac{m(n)}{n}\leq\frac{1}{(m(n)-1)!}\to 0$
as $n\to\infty$. Thus, the point $(L\circ\sigma\circ L)(x)=L(\sigma(y))$ belongs to $B_{k}$ and therefore $\sigma(y)\in L(B_{k})\subset N$ as desired.∎
4.2. The set $N$ has full Hausdorff dimension
Lemma 4.6.
The set $N=\bigcup_{k\in\mathbb{N}}L(B_{k})$ has Hausdorff dimension $\dim_{H}X$.
Proof.
Since $L$ is an isometry, it satisfies $\dim_{H}L(Z)=\dim_{H}Z$ for every $Z\subset X$ (see for example [missing, Proposition 3.3]). Since $\mu_{k}(B_{k})=1$, by definition we have $\dim_{H}\mu_{k}\leq\dim_{H}B_{k}$. As $\alpha_{k}\to P_{00}$, the matrices $Q_{k}$ converge to $P$ and the vectors $\{{\bf q}_{k}\}$ converge to ${\bf p}$. Thus, by continuity of the expression (11), we have that $h(\mu_{k})\to h(m)=\log\lambda$.
Putting together these facts and using Proposition 2.5 and Proposition 3.3 we have
$\dim_{H}X$ $\geq\dim_{H}N=\sup_{k\geq 1}\dim_{H}L(B_{k})=\sup_{k\geq 1}\dim_{H}B_{k}$
$\geq\sup_{k\geq 1}\dim_{H}\mu_{k}=\sup_{k\geq 1}h(\mu_{k})=\log\lambda=\dim_{H}X.\qed$
4.3. The set $N$ has full topological entropy
The argument for topological entropy is almost the same as for Hausdorff dimension. The extra difficulty in this case is that the map $L$ does not necessarily preserves entropy, as it does not commute with the dynamics $\sigma$. However, we can fix this with the following result.
Proposition 4.7.
For every $k\in\mathbb{N}$, we have $h(\sigma|L(B_{k}))\geq h(\mu_{k})$.
In order to prove Proposition 4.7 we need the following lemma.
Lemma 4.8.
Fix $m\in\mathbb{N}$. Then for $\mu_{k}$ almost every $x\in B_{k}$, we have
$\lim_{n\to\infty}-\frac{1}{n}\log\mu_{k}(C_{n+m}(L(x)))=h(\mu_{k}).$
Proof.
Let $\mu_{k},Q_{k},{\bf q}_{k}$ be denoted respectively $\mu,Q,{\bf q}$ to avoid too much notation. Observe that we have the following Birkhoff average-like expression
$-\frac{1}{n}\log\mu(C_{n}(x))$ $=-\frac{1}{n}\log(q_{x_{0}}Q_{x_{0}x_{1}}Q_{x_{1}x_{2}}\cdots Q_{x_{n-2}x_{n-1}})$
$=-\frac{1}{n}\log q_{x_{0}}+\frac{1}{n}\sum_{i=0}^{n-2}-\log Q_{x_{i}x_{i+1}}$
$=-\frac{1}{n}\log q_{x_{0}}+\frac{1}{n}\sum_{i=0}^{n-2}\psi(\sigma^{i}(x)),$
where $\psi(x)=-\log Q_{x_{0}x_{1}}$. We will use the idea from [PesinPitskel] that splits the original expression in two sums. Let $S_{1}(n):=\{0,1,\ldots,n-1\}\cap S_{1}$ and $S_{2}(n):=\{0,1,\ldots,n-1\}\cap S_{2}$. Since $m$ is fixed and negligible for $n$ large, we write
$\lim_{n\to\infty}-\frac{1}{n}\log\mu(C_{n+m}(L(x)))=\lim_{n\to\infty}\frac{1}{n}\left(\sum_{i\in S_{1}(n)}\psi(\sigma^{i}(L(x)))+\sum_{i\in S_{2}(n)}\psi(\sigma^{i}(L(x)))\right).$
Claim: We have $\lim_{n\to\infty}\frac{1}{\lvert S_{1}(n)\rvert}\sum_{i\in S_{1}(n)}\psi(\sigma^{i}(L(x)))=h(\mu)$.
Proof of Claim:.
Since we are considering indices in the set $S_{1}$ where the map $L$ does not change the symbols, the limit on the left hand side equals $\lim_{n\to\infty}\frac{1}{\lvert S_{1}(n)\rvert}\sum_{i\in S_{1}(n)}\psi(\sigma^{i}(x))$. We cannot use the Strong Law of Large Numbers as Pesin and Pitskel do because our variables are not independent. We need to use the following result [J, K], which establishes the SLLN for quasi uncorrelated variables. It will be written in probabilistic notation.
Proposition 4.9 (Janisch, Korchevsky).
Let $(X_{n})_{n}$ be a sequence of non-negative random variables on a probability space and let $S_{n}=X_{1}+\cdots+X_{n}$. Assume they satisfy:
  1. (i)
    each $X_{n}$ has finite variance $V(X_{n})$, and
    $\sup_{n\in\mathbb{N}}\frac{\mathbb{E}(S_{n})}{n}<\infty\quad\text{and}\quad\sum_{n=1}^{\infty}\frac{V(X_{n})}{n^{2}}<\infty;$
  2. (ii)
    there exists a constant $c>0$ such that
    $V(S_{n})\leq c\sum_{k=1}^{n}V(X_{k})$
    for every $n\in\mathbb{N}$.
Then, the SLLN holds, i.e.
$\lim_{n\to\infty}\frac{S_{n}-\mathbb{E}(S_{n})}{n}=0$
almost surely.
We will show that the sequence $X_{n}(x):=-\log Q_{x_{n-2}x_{n-1}}$ for $n\in S_{1}\cup\{0\}$ satisfies assumptions $(i)$ and $(ii)$. Denote by $S^{*}_{n}:=\sum_{k\in S_{1}(n)}X_{k}$.
First observe that since the measure $\mu$ is $\sigma$-invariant and $X_{n}=\psi\circ\sigma^{n}$, we have $\mathbb{E}(X_{n})=\mathbb{E}(X_{0})$ and $V(X_{n})=V(X_{0})$. It is clear now that $(i)$ holds. For $(ii)$, we have
$V(S^{*}_{n})$ $=\sum_{k\in S_{1}(n)}V(X_{k})+2\sum_{\{i,j\in S_{1}(n):i<j\}}\operatorname{Cov}(X_{i},X_{j})$
$=\lvert S_{1}(n)\rvert V(X_{0})+2\sum_{\{i,j\in S_{1}(n):i<j\}}\operatorname{Cov}(X_{i},X_{j}).$
Observe that for $i<j$, since $\sigma$ preserves $\mu$ we have
$\operatorname{Cov}(X_{i},X_{j})$ $=\mathbb{E}(X_{i}X_{j})-\mathbb{E}(X_{i})\mathbb{E}(X_{j})$
[…]
$\leq Ce^{-\gamma(j-i)}$
for some $\gamma>0$, consequence of the exponential decay of correlations of the Markov measure $\mu$. Here $C$ depends on $\psi$, which depends only on the matrix $Q$.
Now write $S_{1}(n):=A_{1}\cup A_{2}\cup\cdots\cup A_{k_{n}}$, where $k_{n}=\max\{k\in\mathbb{N}:(2k)!\leq n\}$, $A_{\ell}=\{(2\ell)!,\ldots,(2\ell+1)!\}$ for $11\leq\ell\leq k_{n}-$ 1 and
$A_{k_{n}}=\left\{\begin{array}[]{lcl}\{(2k_{n})!,\ldots,(2k_{n}+1)!\}&,&(2k_{n}+1)!\leq n\\ \{(2k_{n})!,\ldots,n\}&,&(2k_{n}+1)!>n\end{array}\right..$
Observe that for $1\leq\ell\leq k_{n}$,
(15) $\sum_{\{i,j\in A_{\ell}:i<j\}}e^{-\gamma(j-i)}=\sum_{i=1}^{\lvert A_{\ell}\rvert-1}(\lvert A_{\ell}-i\rvert)e^{-\gamma i}\leq\lvert A_{\ell}\rvert E,$
where $E=\sum_{n=1}^{\infty}e^{-\gamma n}$. Let $C^{\prime}>0$ be a constant with $E\leq C^{\prime}V(X_{0})$. Then
(16) $\sum_{\ell=1}^{k_{n}}\sum_{\{i,j\in A_{\ell}:i<j\}}e^{-\gamma(j-i)}\leq\sum_{\ell=1}^{k_{n}}\lvert A_{\ell}\rvert E\leq C^{\prime}\lvert S_{1}(n)\rvert V(X_{0}).$
For the reminding terms we have
$\sum_{\ell=1}^{k_{n}}\sum_{m=1}^{\ell-1}\sum_{i\in A_{m}}\sum_{j\in A_{\ell}}e^{-\gamma(j-i)}$ $\leq\sum_{\ell=1}^{k_{n}}\sum_{m=1}^{\ell-1}\sum_{i\in A_{m}}\sum_{j\in A_{\ell}}e^{-\gamma((2\ell)!-i)}$
$=\sum_{\ell=1}^{k_{n}}\lvert A_{\ell}\rvert\underbrace{\sum_{m=1}^{\ell-1}\sum_{i\in A_{m}}e^{-\gamma((2\ell)!-i)}}_{\leq E}$
$\leq\lvert S_{1}(n)\rvert E$
$\leq C^{\prime}\lvert S_{1}(n)\rvert V(X_{0}).$
Putting everything together, we obtain
$V(S_{n}^{*})$ $\leq\lvert S_{1}(n)\rvert V(X_{0})+4CC^{\prime}\lvert S_{1}(n)\rvert V(X_{0})\leq(1+4CC^{\prime})\lvert S_{1}(n)\rvert V(X_{0}).$
Thus, our sequence of random variables satisfies $(ii)$ as well, and hence we conclude the SLLN, i.e. for $\mu$-almost every $x\in B_{k}$ we have
[…]
$=\sum_{w\in\mathcal{L}_{2}(X)}-\mu([w])\log Q_{w}$
$=\sum_{w\in\mathcal{L}_{2}(X)}-q_{w_{0}}Q_{w_{0}w_{1}}\log Q_{w_{0}w_{1}}$
$=h(\mu).$
This completes the proof of the claim. ∎
In order to complete the proof of Lemma 4.8 let $\iota\colon X\to X$ be the involution $0\mapsto 1$, $1\mapsto 0$. Then, $\iota_{*}\mu$ is a Markov measure with the same entropy as $\mu$. Using an analogous argument as for $\mu$, we obtain that for $\mu$-almost every $x\in B_{k}$
$\lim_{n\to\infty}\frac{1}{\lvert S_{2}(n)\rvert}\sum_{i\in S_{2}(n)}\psi(\sigma^{i}(L(x)))$ $=\lim_{n\to\infty}\frac{1}{\lvert S_{2}(n)\rvert}\sum_{i\in S_{2}(n)}\psi(\sigma^{i}(\iota(x)))$
$=\lim_{n\to\infty}\frac{1}{\lvert S_{2}(n)\rvert}\sum_{i\in S_{2}(n)}\psi(\iota(\sigma^{i}(x)))$
$=h(\iota_{*}\mu)=h(\mu).$
Finally since $\lvert S_{1}(n)\rvert+\lvert S_{2}(n)\rvert=n$, we have
$\lim_{n\to\infty}-\frac{1}{n}\log\mu(C_{n+m}(L(x)))$ $=\lim_{n\to\infty}\frac{1}{n}\left(\sum_{i\in S_{1}(n)}\psi(\sigma^{i}(L(x)))+\sum_{i\in S_{2}(n)}\psi(\sigma^{i}(L(x)))\right)$
$=\lim_{n\to\infty}\left(\frac{\lvert S_{1}(n)\rvert}{n}\frac{1}{\lvert S_{1}(n)\rvert}\sum_{i\in S_{1}(n)}\psi(\sigma^{i}(L(x)))\right.$
$+\left.\frac{\lvert S_{2}(n)\rvert}{n}\frac{1}{\lvert S_{2}(n)\rvert}\sum_{i\in S_{2}(n)}\psi(\sigma^{i}(L(x)))\right)$
$=h(\mu)$
as desired.∎
Proof of Proposition 4.7.
We follow the line of arguments as in [PesinPitskel]. We are going to use the definition of entropy and notation from subsection 2.3. Let $\delta>0$ and fix $m\in\mathbb{N}$. Consider the open cover $\mathcal{U}$ of $X$ by cylinders of length $m$. Observe that for ${\bf U}\in\mathcal{S}$, the set $X({\bf U})$ is a cylinder of length $m+m({\bf U})$.
By Lemma 4.8, given $0<\delta<\min\{h(\mu_{k}),1/2\}$ there is $D\subset B_{k}$ and $N\in\mathbb{N}$ such that $\mu_{k}(D)\geq 1-\delta$ and for every $x\in D$ we have
(17) $\mu(C_{n+m}(L(x)))\leq e^{-n(h(\mu_{k})-\delta)}$
for every $n\geq N$.
Let $\alpha<h(\mu_{k})-\delta$ and denote the Carathéodory outer measure by $H:=m_{C}(\cdot,\alpha)$. Let $n\geq N$ and choose $\mathcal{G}\subset\mathcal{S}$ a collection of strings that cover $L(B_{k})$ with $m({\bf U})\geq n$ for every ${\bf U}\in\mathcal{G}$ such that
(18) $\Big{|}\inf\left\{\sum_{{\bf V}\in\mathcal{J}}e^{-m({\bf V})\alpha}\right\}-\sum_{{\bf U}\in\mathcal{G}}e^{-m({\bf U})\alpha}\Big{|}\leq\delta,$
where the infimum is taken over all $\mathcal{J}\subset\mathcal{S}$ that cover $L(B_{k})$ with $m({\bf V})\geq n$ for every ${\bf V}\in\mathcal{J}$. Let $\mathcal{G}_{r}:=\{{\bf U}\in\mathcal{G}:m({\bf U})=r\}$ and $E_{r}:=\bigcup_{{\bf U}\in\mathcal{G}_{r}}X({\bf U})$. Since $X({\bf U}^{\prime})$ and $X({\bf U}^{\prime\prime})$ are disjoint for distinct ${\bf U}^{\prime},{\bf U}^{\prime\prime}$ by (17) we have
(19) $\mu_{k}(L(E_{r})\cap D)=\sum_{{\bf U}\in\mathcal{G}_{r}}\mu_{k}(L(X({\bf U}))\cap D)\leq\lvert\mathcal{G}_{r}\rvert e^{-r(h(\mu_{k})-\delta)}.$
Then,
$\sum_{{\bf U}\in\mathcal{G}}e^{-m({\bf U})\alpha}$ $=\sum_{r=n}^{\infty}e^{-r\alpha}\lvert\mathcal{G}_{r}\rvert$
$\geq\sum_{r=n}^{\infty}\mu_{k}(L(E_{r})\cap D)e^{r(-\alpha+h(\mu_{k})-\delta)}$
$\geq(1-\delta)e^{n(-\alpha+h(\mu_{k})-\delta)}$
$\geq 1-\delta$
as $-\alpha+h(\mu_{k})-\delta>0$.
It follows that
$\inf\left\{\sum_{{\bf V}\in\mathcal{J}}e^{-m({\bf V})\alpha}:\mathcal{J}\text{ covers }L(B_{k}),m({\bf V})\geq n\right\}\geq 1-2\delta>0.$
Taking limit as $n\to\infty$ we obtain $H(L(B_{k}))=m_{C}(L(B_{k}),\alpha)>0$. By definition, we get $h_{\mathcal{U}}(\sigma|L(B_{k}))\geq\alpha$. Since $\alpha<h(\mu_{k})-\delta$ is arbitrary, we obtain $h_{\mathcal{U}}(\sigma|L(B_{k}))\geq h(\mu_{k})-\delta$. Now since $\delta$ is arbitrarily close to 0, we get $h_{\mathcal{U}}(\sigma|L(B_{k}))\geq h(\mu_{k})$. Finally we can send the length of the cylinders in the cover $\mathcal{U}$ to infinity, and observing that $\operatorname{diam}\mathcal{U}\to 0$ as $m\to\infty$ we get the desired result.∎
We now proceed to show that $N=\bigcup_{k\in\mathbb{N}}L(B_{k})$ has full topological entropy. By Proposition 2.7 and Proposition 4.7, we have
$h_{top}(\sigma)$ $\geq h(\sigma|N)=\sup_{k\in\mathbb{N}}h(\sigma|L(B_{k}))$
$\geq\sup_{k\in\mathbb{N}}h(\mu_{k})=\log\lambda$
$=h_{top}(\sigma)$
as desired. ∎
4.4. Finding an uncountable collection $\mathcal{N}$ of such sets $N$.
Let $A\subset(0,\infty)$ be an uncountable set of points which are pairwise rationally independent, i.e. for every $x,y\in A$ if $k_{1}x+k_{2}y=0$ with $k_{1},k_{2}\in\mathbb{Z}$, then $k_{1}=k_{2}=0$. For every $a\in A$, consider the sequence
(20) $\mathcal{P}(a):=\{P_{00}-\frac{a}{m}:m\geq m_{0}\}=\{\alpha_{k}(a):k\in\mathbb{N}\},$
starting from $m$ such that all the terms of the sequence are positive and all the matrices $Q_{k}$ defined by this sequence are aperiodic. This sequence gives us one of the sets we constructed $N_{\mathcal{P}(a)}$.
We claim that for $a\neq b$ we have $\mathcal{P}(a)\cap\mathcal{P}(b)=\emptyset$. Indeed, if they have intersection, then there are integer numbers $m,m^{\prime}$ such that
$P_{00}-\frac{a}{m}=P_{00}-\frac{b}{m^{\prime}},$
therefore $m^{\prime}a-mb=0$ which contradicts rationally independence of $a$ and $b$. Now observe that if $\mathcal{P}(a)\cap\mathcal{P}(b)=\emptyset$, then
$\left(\bigcup_{k\in\mathbb{N}}B_{k}^{\mathcal{P}(a)}\right)\cap\left(\bigcup_{k\in\mathbb{N}}B_{k}^{\mathcal{P}(b)}\right)=\emptyset,$
as these sets are defined such that their points have Birkhoff averages the elements of each sequence. Since $L\colon X\to X$ is a bijection, we conclude that $N_{\mathcal{P}(a)}\cap N_{\mathcal{P}(b)}=\emptyset$ whenever $a\neq b$.
Thus, the collection
$\mathcal{N}:=\{N_{\mathcal{P}(a)}:a\in A\}$
is an uncountable collection of subsets of $X$ satisfying properties (1)–(4). This completes the proof of Theorem 3.6. ∎
5. Extending to a larger class of examples
In this section we provide an example of a topologcally mixing subshift of finite type that does not satisfy the property of having two safe symbols, however one can still prove the result. The assumption is required so that the map $L$ is well defined.
Let $A=\begin{pmatrix}1&1\\ 1&0\end{pmatrix}$ and let
$X:=\{(x_{i})_{i\geq 0}:x_{i}\in\{0,1\},A_{x_{i}x_{i+1}}=1\text{ for every }i\geq 0\}.$
The system $(X,\sigma)$ is topologically mixing and does not have two different safe symbols. Observe that $X$ is obtained by forbidding the word $11$ in the full shift on two symbols. Consider a new alphabet $\mathcal{A}$ given by admissible words of length 3 in $X$. That is,
$\mathcal{A}=\{000,010,001,100,101\}.$
Denote the symbols by $a=000,b=010,c=001,d=100,e=101$. Consider the adjacency matrix $A^{\prime}$ given by this alphabet and forbidding the word 11. Then,
$A^{\prime}=\begin{pmatrix}1&1&1&1&1\\ 1&1&1&1&1\\ 1&1&1&0&0\\ 1&1&1&1&1\\ 1&1&1&0&0\end{pmatrix}.$
Let $Y\subset\{a,b,c,d,e\}^{\mathbb{N}_{0}}$ be the subshift of finite type defined by the adjacency matrix $A^{\prime}$, and observe that $a$ and $b$ are safe symbols. Let $\pi\colon X\to Y$ be defined by $\pi((x_{i})_{i\geq 0})=(y_{i})_{i\geq 0}$, where $y_{i}=x_{3i}x_{3i+1}x_{3i+2}$. Briefly, $\pi$ sends a sequence $x\in X$ to $Y$ by “grouping” symbols of length 3:
$\pi(x_{0}x_{1}x_{2}x_{3}x_{4}x_{5}x_{6}x_{7}x_{8}\ldots)=(x_{0}x_{1}x_{2})(x_{3}x_{4}x_{5})(x_{6}x_{7}x_{8})\ldots\in Y.$
Proposition 5.1.
Using the metric (9) on both $X$ and $Y$, the map $\pi$ is a bi-Hölder homeomorphism.
Proof.
Let $x,x^{\prime}\in X$ be two distinct sequences such that the first index where they differ is $3k+j$ for some $k\in\mathbb{N}$ and $j\in\{0,1,2\}$. Then, $d(x,x^{\prime})=e^{-(3k+j)}$. Observe that the first symbol where $\pi(x)$ and $\pi(x^{\prime})$ differ is $k$. Since
$(e^{-(3k+j)})^{1/3}\leq e^{-k}=e^{j/3}(e^{-(3k+j)})^{1/3}\leq e(e^{-(3k+j)})^{1/3},$
we get
(21) $d(x,x^{\prime})^{1/3}\leq d(\pi(x),\pi(x^{\prime}))\leq ed(x,x^{\prime})^{1/3}$
as desired.∎
Moreover, $\pi$ is a topological conjugacy between $(Y,\sigma_{Y})$ and $(X,\sigma^{3})$. As a consequence $(Y,\sigma_{Y})$ is also topologically mixing. The idea now is to use our result in $(Y,\sigma_{Y})$ and transfer the result to the system $(X,\sigma)$. However, we must be careful as irregular points in $(Y,\sigma_{Y})$ will be sent to irregular points in $(X,\sigma^{3})$, and those points are not necessarily irregular for $(X,\sigma)$.
Theorem 5.2.
Let $X$ be as in our previous example. Then, there is an uncountable collection $\mathcal{N}$ of subsets of $X$ satisfying:
  1. (1)
    There is $\varphi\in C(X)$ such that every $N\in\mathcal{N}$ is contained in $\hat{X}(\varphi)$;
  2. (2)
    every $N\in\mathcal{N}$ is dense in $X$ and satisfies $\sigma^{3}(N)\subset N$;
  3. (3)
    for every distinct $N,N^{\prime}\in\mathcal{N}$ we have $N\cap N^{\prime}=\emptyset$;
  4. (4)
    for every $N\in\mathcal{N}$ we have $\dim_{H}N=h(\sigma|N)=\log\left(\frac{1+\sqrt{5}}{2}\right)$.
Proof.
We will use the following result ([GPS]). We specify now the dynamical systems with respect to which we are considering irregular points.
Proposition 5.3.
For every $\phi\in C(X)$, $\hat{X}(\sigma^{3},\phi+\phi\circ\sigma+\phi\circ\sigma^{2})=\hat{X}(\sigma,\phi)$.
We now replicate the proof of Theorem 3.6 on $(Y,\sigma)$ with a slightly different function. Let $\psi,g\colon Y\to\mathbb{R}$ be defined by
(22) $\psi(y):=\left\{\begin{array}[]{lll}3&,&y_{0}=a\\ 2&,&y_{0}\in\{b,c,d\}\\ 1&,&y_{0}=e\end{array}\right.,\qquad g(y):=\left\{\begin{array}[]{lll}3&,&y_{0}=b\\ 2&,&y_{0}\in\{a,c,d\}\\ 1&,&y_{0}=e\end{array}\right.$
Let $\mu$ be a Markov measure on $Y$ with matrix $Q$ and probability vector ${\bf q}$ such that $q_{a}\neq q_{b}$. Observe that
[…]
Let $B(\mu)\subset Y$ be the set of $y\in Y$ such that
[…]
Define the function $L\colon Y\to Y$ as follows: for $y=(y_{i})_{i\geq 0}$,
  • if $i\in S_{1}$, set $L(y)_{i}=y_{i}$;
  • if $i\in S_{2}$, set
    $L(y)_{i}=\left\{\begin{array}[]{ccl}a&\text{if}&y_{i}=b\\ b&\text{if}&y_{i}=a\\ y_{i}&\text{if}&y_{i}\in\{c,d,e\}\end{array}\right..$
Fix $y\in B(\mu)$, and observe that if $(2n-1)!+1\leq j\leq(2n)!-1$ for some $n\geq 0$ we have
$\psi(\sigma^{j}(L(y)))=\left\{\begin{array}[]{ccl}3&\text{if}&L(y)_{j}=a\\ 2&\text{if}&L(y)_{j}\in\{b,c,d\}\\ 1&\text{if}&L(y)_{j}=e\end{array}\right.=\left\{\begin{array}[]{ccl}3&\text{if}&y_{j}=b\\ 2&\text{if}&y_{j}\in\{a,c,d\}\\ 1&\text{if}&y_{j}=e\end{array}\right.=g(\sigma^{j}(y)).$
[…]
Then, there is an uncountable collection $\mathcal{R}$ of subsets of $Y$ satisfying:
  1. (1)
    every $R\in\mathcal{R}$ is contained in $\hat{Y}(\sigma,\psi)$;
  2. (2)
    every $R\in\mathcal{R}$ is dense in $Y$ and satisfies $\sigma_{Y}(R)\subset R$;
  3. (3)
    for every distinct $R,R^{\prime}\in\mathcal{R}$ we have $R\cap R^{\prime}=\emptyset$;
  4. (4)
    for every $R\in\mathcal{R}$ we have $\dim_{H}R=h(\sigma_{Y}|R)=\dim_{H}Y=h_{top}(Y,\sigma_{Y})$.
Consider $\varphi(x):=\left\{\begin{array}[]{lll}1&,&x_{0}=0\\ 0&,&x_{0}\neq 0\end{array}\right..$
Proposition 5.4.
The collection $\mathcal{N}:=\{\pi^{-1}(R):R\in\mathcal{R}\}$ satisfies:
  1. (1)
    every $\pi^{-1}(R)\in\mathcal{N}$ is contained in $\hat{X}(\sigma,\varphi)$;
  2. (2)
    every $\pi^{-1}(R)\in\mathcal{N}$ is dense in $X$ and satisfies $\sigma^{3}(\pi^{-1}(R))\subset\pi^{-1}(R)$;
  3. (3)
    for every distinct $\pi^{-1}(R),\pi^{-1}(R^{\prime})\in\mathcal{N}$ we have $\pi^{-1}(R)\cap\pi^{-1}(R^{\prime})=\emptyset$;
  4. (4)
    for every $\pi^{-1}(R)\in\mathcal{N}$ we have $\dim_{H}\pi^{-1}(R)=\dim_{H}X$ and $h(\sigma|\pi^{-1}(R))=h_{top}(X,\sigma)$
Proof.
Property (3) and density of $\pi^{-1}(R)$ follow from the corresponding property for $\mathcal{R}$ and the fact that $\pi$ is a homeomorphism.
For (1), let $\eta\colon X\to\mathbb{R}$ be defined by $\eta:=\varphi+\varphi\circ\sigma+\varphi\circ\sigma^{2}$. Observe that $\eta=\psi\circ\pi$. Let $x\in\pi^{-1}(R)\in\mathcal{N}$. Then since $\pi$ is a conjugacy between $(X,\sigma^{3})$ and $(Y,\sigma_{Y})$, we have $x\in\hat{X}(\sigma^{3},\eta)=\hat{X}(\sigma,\varphi)$ by Proposition 5.3.
For (4), by (21) we have that $\pi$ is Hölder with exponent $1/3$ and $\pi^{-1}$ is Hölder with exponent $3$. By [missing, Proposition 2.3] for every $Z\subset X$ we have $\dim_{H}\pi(Z)=3\dim_{H}Z$. Let $R\in\mathcal{R}$, then
$\dim_{H}\pi^{-1}(R)$ $=\dfrac{1}{3}\dim_{H}R=\frac{1}{3}\dim_{H}Y$
$=\frac{h_{top}(Y,\sigma_{Y})}{3}=\dfrac{h_{top}(X,\sigma^{3})}{3}$
$=h_{top}(X,\sigma)=\dim_{H}X.$
Similarly,
$h(\sigma|\pi^{-1}(R))$ $=\frac{h(\sigma^{3}|\pi^{-1}(R))}{3}=\dfrac{h(\sigma_{Y}|R)}{3}$
$=\frac{h_{top}(Y,\sigma_{Y})}{3}=\dfrac{h_{top}(X,\sigma^{3})}{3}$
$=h_{top}(X,\sigma).$
Proposition 5.4 finishes the proof of Theorem 5.2. ∎
This example suggest that topologically mixing subshifts of finite type that allows (up to a power of the dynamics) a conjugacy to a system $(Y,\sigma_{Y})$ given by grouping symbols into words of larger length may still have the chance to satisfy a version of the conclusion to our theorem, even if they do not satisfy the safe symbols condition. This example also shows that the safe symbols condition is a sufficient condition to have the conclusion to our theorem, but not a necessary one.